### v3

# Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs

Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in language-related tasks. However, their deployment presents significant challenges due to their substantial memory and storage requirements. To address this challenge, weight-only quantization has emerged as a promising solution. Previous research has indicated that fine-tuning through up and down rounding can enhance performance. In this study, we introduce SignRound, a method that utilizes signed gradient descent (SignSGD) to optimize rounding values and weight clipping within just 200 steps, combining the strengths of both Quantization-Aware Training (QAT) and Post-Training Quantization (PTQ). SignRound achieves outstanding results compared to recent methods across 2 to 4 bits, while maintaining low tuning costs and without introducing any additional inference overhead. For instance, SignRound led to absolute average accuracy improvements ranging from 6.91% to 33.22% at 2 bits. Furthermore, it demonstrates robust generalization to various recent models and achieves near-lossless quantization in most scenarios at 4 bits. The source code is publicly available at https://github.com/intel/auto-round.
1 Introduction
Large language models (LLMs) have demonstrated remarkable proficiency in language-related tasks (OpenAI, 2022; Touvron et al., 2023a). However, deploying LLMs poses significant challenges due to their extensive memory and storage requirements. Additionally, the computational demands of these models create obstacles for real-time applications. Thus, studying techniques such as quantization becomes crucial to enable the efficient deployment of LLMs. Quantization techniques can be broadly categorized into two main types: quantization-aware training (QAT) (Esser et al., 2019; Zhuang et al., 2021; Lee et al., 2021; Liu et al., 2023b) and post-training quantization (PTQ) (Nagel et al., 2019; Xiao et al., 2022; Frantar et al., 2022; Nagel et al., 2020).
QAT involves training the model with quantization in mind. During QAT, the model is trained using simulated lower-precision representations, allowing it to learn and adapt to the effects of quantization. This approach often results in better accuracy compared to PTQ. However, QAT does come with certain drawbacks, including increased training complexity, longer training times, and the need to tune hyperparameters. The application of QAT to LLMs can be particularly costly, despite recent efforts (Hu et al., 2021; Dettmers et al., 2023) to improve the efficiency of fine-tuning LLMs.
On the other hand, PTQ directly quantizes the model without any simulated training or fine-tuning. While PTQ is a more straightforward approach, it is susceptible to significant accuracy drops. This underscores the importance of further advancements in PTQ methods to enhance their accuracy preservation capabilities.
Two types of tensors commonly undergo quantization: activations and weights. Weight-only quantization has recently gained traction for LLMs due to its advantageous tradeoff. Quantizing activations for LLMs can be complex (Wei et al., 2022b; Xiao et al., 2023; Bondarenko et al., 2023), making weight-only quantization a more practical and feasible option. Moreover, the main bottleneck in generating new tokens for LLMs often arises from memory bandwidth limitations (Kim et al., 2023a), emphasizing the relevance of weight-only quantization.
This study focuses on weight-only quantization. In quantizing weights, a critical step involves rounding, primarily achieved through rounding-to-nearest (RTN). RTN quantizes each weight independently by rounding it to the nearest integer. However, RTN overlooks the relationships between weights as well as between weights and activations. Nagel et al. (Nagel et al., 2020) explored the potential for an enhanced rounding strategy to improve accuracy. They approached the rounding task by formulating it as a quadratic unconstrained binary optimization problem and approximating the loss using a Taylor series expansion. However, relying solely on the second-order term may not yield accurate results, as rounding can significantly modify weights, making other order terms non-negligible.
We chose SignSGD as our optimization method to approach the optimal rounding solution within a limited number of steps, drawing inspiration from the well-defined boundaries of the solution space. This space is confined to ranges of [-0.5, 0.5] for rounding and [0, 1] for weight clip scales, offering several advantages for SignSGD. Firstly, the optimal value for up and down rounding typically resides in a large region rather than a single float, as only the threshold for altering the rounding value is significant. This eliminates the necessity for the gradient magnitude to converge precisely to a single point. Secondly, due to these confined boundaries, SignSGD allows efficient navigation of this space within a limited number of steps. In contrast, optimizers like Adam (Kingma and Ba, 2014) may struggle due to significant variations in gradient magnitude, making it challenging to converge to the optimal value within a restricted number of steps. Thirdly, SignSGD is inherently intuitive, facilitating easy adjustment of the step size (learning rate). For example, we employed the same optimizer hyperparameters across all experiments unless explicitly stated, consisting of 200 steps and a learning rate of $35e-$ 3, with linear weight decay. This ensures that $200*0.005/2=0.5$ covers the range of [-0.5, 0.5] for rounding and [0.5, 1] for weight clipping, which works well in practice. Fourthly, SignSGD stands out for its lightweight nature compared to other optimizers, requiring less memory and computational resources. Figure 1 provides an overview of our method. Our contributions primarily lie in three aspects:
  • We introduce a concise yet effective method for optimizing the weight only quantization, combining the strengths of both QAT and PTQ. Our approach leverages SignSGD to tune the rounding with the weight clip, without introducing any additional overhead during inference.
  • Our empirical results demonstrate a notable or significant enhancement in performance compared to recent works across different quantization configurations, from 2-bit to 4-bit.
  • We demonstrate that SignRound’s performance can be further enhanced by fine-tuning model-wise hyperparameters within a constrained space. Moreover, our method demonstrates strong generalization across various models and delivers nearly lossless results across the majority of scenarios using 4-bit quantization.
2 Related Work
Quantization Aware Training.
QAT methods have gained widespread popularity in model compression, as they enable the fine-tuning process  (Esser et al., 2019; Zhuang et al., 2021; Lee et al., 2021)  (Zhuang et al., 2021), often leading to superior accuracy compared to the PTQ method,
Post-training Quantization (PTQ).
PTQ methods simplify the quantization process without the needs of additional training. (Nagel et al., 2019; Liu et al., 2021; Frantar and Alistarh, 2022; Hassibi et al., 1993; Yao et al., 2021; Wang et al., 2019). Given its low resource requirement, PTQ is particularly suitable for the quantization of Large Language Models (LLMs).
Large Language Models Quantization.
Significant strides have been made in addressing the pressing need for quantizing large language models (LLMs). LLM.int8() (Dettmers et al., 2022) introduces a mixed-precision approach to preserve crucial channels in high precision. PEQA (Kim et al., 2024) presents a simple yet effective method that combines the benefits of parameter-efficient fine-tuning (PEFT) with quantized LLMs. AQLM (Mao et al., 2024) builds upon Additive Quantization, a classic algorithm from the Multi-Codebook Quantization family, adapting it to LLM quantization. ZeroQuantV2 (Yao et al., 2023) employs low-rank matrices to enhance model quality recovery. RPTQ (Yuan et al., 2023) addresses range differences between channels by rearranging and quantizing them in clusters. Other methods, such as SPIQ (Yvinec et al., 2023b), SmoothQuant (Xiao et al., 2022), and Outlier Suppression+ (Wei et al., 2023), utilize handcrafted equivalent transformations to mitigate quantization errors. LLM-QAT (Liu et al., 2023b) employs QAT to enhance performance.
Weight only quantization.
Weight-only quantization reduces the memory footprint and bandwidth demands by quantizing only the weights while retaining activations in floating-point precision, offering a promising balance between accuracy and compression. GPTQ (Frantar et al., 2022) optimizes weights using the Optimal Brain Surgeon technique (Hassibi et al., 1993), achieving low-bit quantization on LLMs with minimal tuning overhead. AWQ (Lin et al., 2023) follows the equivalent transformation approach with additional tuning in a constrained space, sharing similar limitations with SmoothQuant (Xiao et al., 2022). TEQ(Cheng et al., 2023) and Omniquant (Shao et al., 2023) both utilize a trainable equivalent transformation, while Omniquant employs extra weight clip tuning. HQQ (Badri and Shaji, 2023) accelerates quantization for large models by eliminating the need for calibration data, making the quantization process extremely fast. Some other works have incorporated optimization methods with extra inference overhead to improve quantization accuracy, such as dense-and-sparse decomposition techniques in SqueezeLLM (Kim et al., 2023a) and EasyQuant (Tang et al., 2024), as well as nonuniform quantization methods in NUPES (Yvinec et al., 2023a), QuIP#(Tseng et al., 2024),(Gong et al., 2024), AQLM(Mao et al., 2024), etc. Additionally, FineQuant (Kim et al., 2023b) introduces a straightforward heuristic weight quantization approach that adaptively determines quantization granularity. In this work, we focus on approaches that do not introduce overhead during inference.
Rounding Methods.
Adaptive Rounding (Nagel et al., 2020) has already showcased the potential of an advanced rounding strategy to enhance accuracy (Li et al., 2021; Wei et al., 2022a). They used the rounding task as a quadratic unconstrained binary optimization problem by approximating the task loss through a Taylor series expansion. However, considering only the second-order term may not yield accurate results. This is because the rounding value gets multiplied by a scaling coefficient during de-quantization, potentially introducing significant weight changes that make other order terms non-negligible. FlexRound (Lee et al., 2023) introduces a more flexible approach to rounding by incorporating element-wise division. However, it’s not easily scalable to apply to LLMs due to the needs of specialized hyperparameters for each specific model and task. Furthermore, Oscillation-free (Liu et al., 2023a) suggests that the introduction of learnable parameters might result in weight oscillation problems. AQuant (Li et al., 2022) introduced a dynamic approach where the border becomes a function dependent on the activation value to reduce the quantization error of activation.
Signed Gradient Descent.
Signed gradient descent is not commonly utilized and is typically applied in specific scenarios, such as reducing communication costs. This is because signed gradient carries significantly less information compared to original gradient. Recent studies have shed light on the advantages of sign-based methods over gradient descent in certain conditions. Safaryan et al. (Safaryan and Richtárik, 2021) found that sign-based methods are preferable when the Hessian matrix is concentrated on its diagonal and the maximal eigenvalue is much larger than the average eigenvalue. Li et al. (Li et al., 2023a) investigated a variant of sign-based gradient descent that exhibits faster convergence. Additionally, Safaryan et al. (Safaryan and Richtárik, 2021) proposed a stochastic sign descent with momentum, which converges under the standard bounded variance assumption with the optimal asymptotic rate. These findings contribute to a better understanding of the potential benefits and applications of signed gradient descent methods.
Figure 1: An illustration of SignRound. Unlike the direct rounding in RTN, SignRound performs signed gradient descent to fine-tune the up and down rounding through block-wise output reconstruction. After lightweight forward and backward steps, WINT4 has been well optimized towards the minimal loss, therefore ready for the final inference deployment. Note that Quant and Dequant are two standard operations for quantization and dequantization respectively.
3 Methodology
We begin with an overview of quantization before delving into the specifics of our approach. The following operations can be utilized to quantize and de-quantize the weights $W$:
$\widetilde{W}=s*clip(\left\lfloor\frac{W}{s}+zp\right\rceil,n,m),n,m\in\mathbb{N}$ (1)
[…]
$s=\frac{max(W)-min(W)}{2^{bit}-1}$ (2)
In order to improve the efficacy of the rounding quantization operation, we build upon prior research (Nagel et al., 2020) by introducing a single trainable parameter $V$ to adjust the rounding values.
$\widetilde{W}=s*clip(\left\lfloor\frac{W}{s}+zp+V\right\rceil,n,m),n,m\in\mathbb{N}$ (3)
Additionally, following recent works (Lin et al., 2023; Shao et al., 2023), we introduce two additional trainable parameters, denoted as $\alpha\in[0,1]$ and $\beta\in[0,1]$, to fine-tune the scale of weight clipping. These parameters are incorporated into the equations as follows:
$s=\frac{max(W)*\alpha-min(W)*\beta}{2^{bit}-1}$ (4)
These modifications enable a more adaptable quantization process. We utilize block-wise output reconstruction to train these parameters via optimizer, thus framing the optimization as follows.
$\min_{\alpha,\beta,V}\|WX-\widetilde{W}X\|_{F}^{2}$ (5)
where $X$ is the input of the block and $||\cdot||_{F}$ denotes the Frobenius norm.
Our method distinguishes itself primarily by leveraging SignSGD, with the motivation thoroughly outlined in Section Introduction 1. Figure 1 provides an illustration of our approach. And the pseudocode 1 presents more details of SignRound.
Algorithm 1 SignRound
Input: Calibration Data $\mathcal{D}$, learning rate $lr$, total steps $T$, Model $M$, block module $m_{w}$ with weights $w$, batch size $bs$
Output: $best\_V$, $best\_\alpha$, $best\_\beta$
1:$V\leftarrow 0$, $\alpha\leftarrow 1.0$, $\beta\leftarrow 1.0$ , $best\_l\leftarrow maximum$
2:for $i\leftarrow 0$ to $T$ do
3:     $d\leftarrow\textit{draw $bs$ samples from privious quantized block}$
4:     $x\leftarrow M(d)_{m}$triangleright get the inputs of m
5:     $y_{f}\leftarrow m_{w}(x)$triangleright get the output of original module
6:     $\widetilde{w}\leftarrow qdq(w,\alpha,\beta,V)$triangleright quantize and dequantize w via Eq.3
7:     $y_{q}\leftarrow m_{\widetilde{w}}(x)$triangleright get the output of quantized module
8:     $loss\leftarrow mse(y_{q},y_{f})$triangleright get the loss via Eq.5
9:     $loss.backward()$
10:     if $loss<best\_l$ then
11:         $best\_V,best\_\alpha,best\_\beta\leftarrow V,\alpha,\beta$
12:         $best\_l\leftarrow loss$
13:     end if
14:     update $\alpha$, $\beta$ and V via SignSGD optimizer
15:end for
4 Experiments
In this section, we conduct a comprehensive evaluation of SignRound from various perspectives. Initially, we offer a concise overview of the LLM architectures and tasks encompassed in our assessment. Next, we delve into a detailed comparison between our method and several existing approaches, highlighting the unique advantages of SignRound. Furthermore, we conduct ablation studies to strengthen the efficacy of our choices and explore the sensitivity of hyperparameters. Lastly, we assess the generation ability of our method across various recent models. The tuning cost comparisons are present in Appendix A.
4.1 Experimental Settings
Evaluation and tasks.
We evaluate multiple language tasks to accommodate the task-agnostic setting. Specifically, we present the average accuracy results for 11 zero shot tasks, including HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), LAMBADA (Paperno et al., 2016), TruthfulQA (Lin et al., 2021), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), RTE (Dagan et al., 2010), ARC-Easy, ARC-Challenge (Clark et al., 2018) and MMLU (Hendrycks et al., 2020). We use lm-eval-harness(Gao et al., 2023) for all the above tasks. Furthermore, we complement our evaluation with perplexity (ppl) analysis on Wikitext2 (Merity et al., 2016) PTB (Marcus et al., 1993) and C4 (Raffel et al., 2020), by following the source code  [^1] https://github.com/IST-DASLab/gptq of GPTQ and Wikitext2 (Merity et al., 2016) using lm-eval-harness(Gao et al., 2023). However, We argue that perplexity is notably influenced by outliers,as illustrated in Table 14 for different algorithms. This susceptibility likely arises from the mathematical expression $\mathrm{PPL}(X)=\exp\left(-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})\right)$, where assigning a low probability to even one token can significantly inflate the perplexity score. Consequently, we prioritize the accuracy of the 11 tasks mentioned above as the primary metric, with perplexity data serving as a supplementary reference.
Quantization Configurations.
In accordance with GPTQ (Frantar et al., 2022), our focus lies specifically on weight-only quantization, targeting the linear layers within transformer blocks. Layers such as the embedding layer and typically the last linear layer like ’lm-head’ are excluded from the quantization process. Our evaluation primarily centers on W4, W4G128, and W3G128 configurations, where W4 indicates quantizing weights with 4 bits and G represents finer-grained grouping as described in (Park et al., 2022; Frantar et al., 2022). We adopt asymmetric quantization for these configurations. For calibration and to mitigate overfitting on the WikeText and C4 datasets, we randomly select 512 samples with the same seed from the readily available pile-10k dataset  [^2] https://huggingface.co/datasets/NeelNanda/pile-10k, which comprises the first 10k samples from pile (Gao et al., 2020). We used a sequence length of 2048 for calibration, while for other methods, we adhere to their official settings
Large Language Models.
We compare different algorithms on LLaMA-V1 (Touvron et al., 2023a), LLaMA-V2 (Touvron et al., 2023b) and Mistral-7B-v0.1 (Jiang et al., 2023) as they are commonly used. We cover a wide range of LLM parameters, ranging from 7B to 70B, to ensure comprehensive coverage and analysis.
Table 1: Average accuracies($\uparrow$) across 11 tasks, detailed in Section 4.1, of LLaMA and Mistral models at W2-W4. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, detailed in Section 4.2 , whereas for the 70B model, we only tested a few options.
Config Method Mistral-7B V2-7B V2-13B V2-70B
FP16 63.30 57.98 61.42 66.12
W4G-1 RTN 58.84 55.49 60.46 65.22
HQQ 58.40 46.05 46.82 57.47
Omni 60.52 56.62 60.31 65.80
GPTQ 61.37 56.76 59.79 65.75
AWQ 61.36 57.25 60.58 66.28
Ours 62.33 57.48 61.20 66.27
Ours* 62.64 57.52 61.23 66.27
W4G128 RTN 62.36 56.92 60.65 65.87
HQQ 62.75 57.41 60.65 66.06
Omni 62.18 57.30 60.51 66.02
GPTQ 62.32 56.85 61.00 66.22
AWQ 62.16 57.35 60.91 66.23
Ours 62.62 57.57 60.85 66.39
Ours* 62.87 57.97 60.90 66.41
Config Method Mistral-7B V2-7B V2-13B V2-70B
FP16 63.30 57.98 61.42 66.12
W3G128 RTN 58.20 53.81 58.57 64.08
HQQ 59.33 54.31 58.10 64.80
Omni 58.53 54.72 59.18 65.12
GPTQ 59.91 54.14 59.58 65.08
AWQ 59.96 55.21 58.86 65.12
Ours 60.43 56.68 59.44 65.31
Ours* 60.96 56.68 59.78 65.59
W2G128 RTN 30.52 29.94 33.51 38.14
HQQ 31.41 29.87 35.28 37.42
Omni 32.17 40.74 46.55 51.31
GPTQ 39.61 35.37 42.46 28.47
AWQ 30.06 30.10 32.16 32.23
Ours 52.71 48.64 53.46 61.69
Ours* 53.01 50.34 54.16 61.77
Table 2: Detrailed accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G-1. Our* denotes the highest accuracy achieved among the 8 hyperparameter choices, whereas for the 70B model, we only tested a few options, detailed in 4.2. Appendix C provides more detailed data.
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
Mistral-7B FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
RTN 55.92 66.10 59.01 71.35 80.14 24.85 29.00 79.17 57.76 77.95 45.99 58.84
GPTQ 58.22 73.45 59.47 74.03 80.20 26.93 31.00 81.50 64.98 78.24 47.01 61.37
AWQ 57.20 71.45 59.21 73.64 79.43 25.34 30.40 82.69 68.95 79.25 47.44 61.36
HQQ 52.65 66.58 59.09 70.56 79.60 23.13 27.80 80.03 59.57 77.02 46.33 58.40
Omni 57.52 70.00 60.27 72.93 79.87 23.99 30.80 81.53 63.90 78.54 46.42 60.52
Ours 59.52 73.76 60.75 73.32 80.09 27.17 33.00 82.02 66.07 80.47 49.49 62.33
Ours* 60.00 73.30 60.57 74.35 80.09 27.91 32.20 83.52 67.51 79.92 49.66 62.64
V2-7B FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
RTN 36.87 67.96 55.63 68.51 76.82 26.19 30.60 73.64 58.84 74.07 41.30 55.49
GPTQ 39.66 71.92 55.89 68.03 77.58 25.09 30.20 76.67 62.09 75.55 41.72 56.76
AWQ 40.24 71.20 56.26 69.61 76.93 26.07 32.60 77.31 63.18 75.00 41.30 57.25
HQQ 28.94 43.96 48.43 59.43 71.82 23.62 24.80 52.11 53.79 64.90 34.73 46.05
Omni 39.82 71.45 55.76 67.56 76.88 25.09 30.80 76.15 64.98 74.12 40.19 56.62
Ours 39.97 71.63 56.52 68.43 77.91 25.70 31.60 76.18 65.70 76.01 42.58 57.48
Ours* 40.85 72.75 56.01 67.88 77.86 25.34 31.80 76.39 66.43 75.88 41.55 57.52
V2-13B FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
RTN 50.37 74.35 59.12 71.98 79.00 24.85 33.00 81.77 64.98 79.08 46.59 60.46
GPTQ 51.14 75.37 59.14 72.06 78.02 25.34 32.20 80.46 62.09 77.36 44.54 59.79
AWQ 51.16 75.98 59.51 70.80 78.40 25.21 34.60 78.26 66.79 79.12 46.59 60.58
HQQ 35.92 49.54 46.27 58.01 72.47 23.99 19.80 61.77 51.26 62.84 33.19 46.82
Omni 51.01 75.45 59.48 71.74 78.94 24.60 33.20 77.37 66.07 78.75 46.76 60.31
Ours 52.30 75.96 59.79 72.30 78.84 25.58 34.00 80.15 66.79 79.38 48.12 61.20
Ours* 52.29 76.15 59.73 71.90 78.51 25.21 34.40 80.24 67.51 79.34 48.21 61.23
V2-70B FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
RTN 63.85 77.62 63.38 76.72 81.50 28.89 37.80 83.39 68.23 81.99 54.10 65.22
GPTQ 64.81 79.27 63.86 76.87 81.61 31.46 36.40 82.23 70.04 82.53 54.18 65.75
AWQ 65.08 78.77 64.14 77.11 81.45 30.48 37.20 83.64 72.92 82.49 55.80 66.28
HQQ 56.45 66.74 53.67 73.32 76.50 25.58 33.40 67.95 61.73 72.90 43.94 57.47
Omni 64.40 79.20 63.91 76.95 81.94 31.70 37.60 82.35 69.31 82.24 54.18 65.80
Ours 65.43 79.55 64.47 78.06 82.10 30.60 36.40 83.91 71.12 82.53 54.78 66.27
SignRound Hyper-parameters.
Unless explicated stated, the tuning process involved adjusting each block for 200 steps , with the learning rate(lr) as 1/steps, i.e. $35e-$ 3, a batch size of 8 and linear lr decay. Additionally, we employed automatic mixed precision (AMP) to accelerate the tuning.
4.2 Comparing With Recent Methods
In this section, we compare our methods with those that have already demonstrated remarkable results and impose no additional overhead on our tested models in weight-only quantization for LLMs, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), HQQ (Badri and Shaji, 2023), Omniquant, (Shao et al., 2023) with a naive method RTN.
To ensure fair comparison as much as possible, we have enabled act-order and true-sequential in GPTQ, and also activated static_group in scenarios where $group\_size\neq-1$. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues. For Omniquant(Shao et al., 2023), we adhere to the official settings, which include running for 20 epochs including W2G128 for saving time and disabling ’let’. We conducted calibration tests using sample sizes of 512 and 128, as well as a sample size of 512 with a batch size of 4. Our findings show that using a sample size of 512 typically results in comparable or slight higher performance for models $<=13B$. Therefore, we present the results based on the sample size of 512. For 70B models, due the the NAN loss issue and to reduce the tuning cost of Omniquant, we adopted 128 samples for calibration.
We present the summary results of Mistral-7B and LLAMA V2 in Table 1 , detailed results of W4G-1 in Table 2 Additional detailed results are provided in Appendix C due to space constraints. In summary, our approach demonstrated superior performance compared to GPTQ (Frantar et al., 2022), achieving scores of 30/32, AWQ (Lin et al., 2023) with 27/32, HQQ (Badri and Shaji, 2023) with 15/16, and OmniQuant (Shao et al., 2023) with a score of 29/32 across llamv1/llamav2/mistral-7b on various quantization settings, including W4G-1, W4G128, W3G128, and W2G128. These evaluations were based on the average accuracies of 11 zero-shot tasks.
It’s worthy to noting that, as the bits going down, the advantages of Signround becomes more notable. For example, at W2G128, SignRound could yield absolute average accuracy improvements ranging from 6.91% to 33.22%.
Moreover, we can enhance the performance by optimizing the model’s hyperparameters from a selection of eight choices, denoted as ours*. These include steps (200, 1000), weight clip lr (1.0/steps, 2.0/steps), and the option to either enable or disable quanted inputs which means utilizing the output from the previous quanted block or the previous original block.
4.3 Comparing With Rounding Methods
In this section, we conduct a comparative analysis between SignRound, FlexRound(Lee et al., 2023), and AdaRound(Nagel et al., 2020). Notably, there is no formal official implementation available for FlexRound and AdaRound for LLMs. Hence, we reference the code  [^3] https://openreview.net/forum?id=-tYCaP0phY_ for further details. However, it’s important to highlight that due to the lack of AMP support and other optimizations, the implementation is notably slow, especially when adhering to the official settings, which involve tuning 5000 steps, as presented in Table 9. Hence, our comparison is limited to models of size 13B or smaller. We set the learning rate to 2e-4 for LLaMA-2-7b and Mistral-7B, and 1e-4 for LLaMA-13-7b to align with the official settings as closely as possible.
Table 3: Comparing with other rounding methods, the average accuracies ($\uparrow$) across 11 tasks (detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
Model Method Iter. Mistral-7B V2-7B V2-13B
W4G-1 Flex 200 58.93 56.10 60.06
1000 60.62 56.98 60.29
5000 60.94 57.49 -
Ada 200 58.30 55.06 59.86
1000 58.38 55.05 59.92
Ours 200 62.33 57.48 61.20
200* 62.64 57.52 61.23
W2G128 Flex 200 30.10 30.01 30.66
1000 30.16 31.26 32.29
Ada 200 30.74 30.21 30.36
1000 30.84 30.30 30.02
Ours 200 52.71 48.64 53.46
200* 53.01 50.34 54.16
4.4 Ablation Study
SignSGD versus Adam.
To validate the effectiveness of SignSGD, Table 4 compares it with the Adam optimizer. SignSGD employs a fixed learning rate of $35e-$ 3 throughout all experiments, comprising 200 steps, with linear weight decay. For Adam, we explored initial learning rates ranging from $32.5e-$ 3 to $22e-$ 2. We choose to quantize models of 13B or less with W4G-1 due to the experiment’s cost. SignSGD demonstrated a distinct advantage in average accuracy metrics across 11 tasks, which demonstrate the unique advantage of signed gradient descent in this scenario.
Table 4: Adam optimizer with various learning rates vs. SignSGD optimizer. The average accuracies($\uparrow$) across 11 tasks (detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
Config Model 2.5e-3 5e-3 7.5e-3 1e-2 1.25e-2 1.5e-2 1.75e-2 2e-2 SignSGD
W4G-1 Mistral-7B 61.82 61.16 61.30 60.69 60.80 61.07 61.53 61.23 62.33
V2-7B 56.79 57.45 57.09 57.28 56.88 57.24 57.40 57.10 57.48
V2-13B 60.58 60.73 60.76 60.86 61.02 60.79 61.06 60.85 61.20
W2G128 Mistral-7B 37.12 40.37 41.11 42.02 42.86 43.55 43.44 42.44 52.71
V2-7B 42.26 44.64 45.08 45.04 45.15 43.13 38.71 35.73 48.64
V2-13B 47.81 50.01 49.55 50.80 48.67 51.94 38.28 34.67 53.46
Round and Weigh clip tuning
To validate the contributions of rounding tuning and weight clip tuning, we conducted ablation studies on three models with two quantization configurations. As shown in Table 5, each component provides benefits over RTN, with rounding tuning offering greater advantages. However, when combined, weight clip tuning can sometimes result in lower accuracy in some cases at W2G128.
Table 5: Ablation study of Round tuning and weight clip tuning. The average accuracies($\uparrow$) across 11 tasks(detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
Config Mistral-7B V2-7B V2-13B Mistral-7B V2-7B V2-13B
W4G-1 W2G128
RTN 58.84 55.49 60.46 30.52 29.94 33.51
Weight clip only 61.10 57.41 60.10 46.60 40.53 49.77
Rounding only 61.62 56.74 60.64 52.32 49.14 54.41
Default 62.33 57.48 61.20 52.71 48.64 53.46
Hypeparameters sensitivity
To validate the sensitivity of hyperparameters in SignRound, we conducted ablation studies on sequence length for calibration, number of samples for calibration, tuning batch size, tuning steps, and tuning learning rate. The results are presented in Table 6. Overall, our default hyperparameters achieved balanced results.
Table 6: Ablation study of of hyperparameter sensitivity. The average accuracies($\uparrow$) across 11 tasks(detailed in 4.1) for LLaMA models at W4G-1.
Model Seqlen_512 Samples_128 Batch_4 Iters_100 LR_1e-2 Default
Mistral-7B 60.32 61.82 61.78 61.06 61.27 62.33
V2-7B 57.91 56.41 57.21 57.10 55.89 57.48
V2-13B 60.88 60.87 61.21 60.80 61.03 61.20
4.5 Generalization to other models
To assess the generalization of our method on LLMs, we evaluate SignRound on various mainstream LLMs like Gemma (Team et al., 2024), Phi (Li et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Llama3 (Touvron et al., 2024). Table 7 demonstrated that all int4 models met the requirement of maintaining accuracy drop within 1% of the FP16 or BF16 accuracy by employing 1000 tuning steps and model wise hype-parameters among 4 choices detailed in scection 4.1. Additionally, one-third of the models exhibited a slight improvement in accuracy compared to the baseline. Please note that the generalization experiments utilize an updated version (0.4.0+) of lm-eval-harness (Gao et al., 2023) and real quantized model, which may result in minor discrepancies compared to other benchmark data.
Table 7: Accuracies($\uparrow$) across 11 tasks(0-shot) with 1000 steps for LLMs at W4G128
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg. Vari. %
gemma-7b FP16 61.62 67.51 60.47 73.24 79.43 30.97 33.20 82.78 65.34 81.78 49.91 62.39 -
Ours 61.47 72.04 59.03 75.14 79.49 30.11 34.00 82.69 70.76 79.59 49.40 63.07 1.09%
gemma-7b-it FP16 50.29 60.35 56.20 67.96 77.09 30.48 37.40 81.38 78.70 75.25 47.27 60.22 -
Ours 49.93 62.86 55.64 67.88 77.31 30.35 37.00 81.44 78.70 75.08 45.73 60.17 -0.08%
Llama-2-7b-chat FP16 46.40 71.05 57.80 66.38 76.39 30.23 33.40 79.76 69.68 73.82 44.20 59.01 -
Ours 45.45 70.37 57.06 66.14 76.33 30.35 32.60 80.64 72.92 73.36 43.52 58.97 -0.07%
Llama-3-8B-Instruct BF16 63.86 71.82 57.69 71.43 78.67 36.23 34.00 82.97 67.51 81.52 52.99 63.52 -
Ours 63.06 72.00 56.99 72.38 77.97 35.37 33.00 83.09 68.59 80.89 51.02 63.12 -0.63%
Mistral-7B BF16 59.61 75.61 61.28 74.43 80.79 28.03 32.80 83.73 66.43 80.85 50.09 63.06 -
Ours 58.80 75.51 60.79 74.51 80.14 28.89 33.00 82.78 69.68 80.60 49.15 63.08 0.03%
Mixtral-8x7B BF16 68.02 78.27 64.90 76.48 82.48 34.27 35.40 85.23 70.76 84.30 56.66 66.98 -
Ours 66.93 78.25 64.59 75.14 82.10 32.19 35.60 84.74 69.31 84.30 56.48 66.33 -0.97%
Mixtral-8x7B-Instruct BF16 68.85 77.18 67.67 76.87 83.51 49.69 36.80 88.50 71.84 86.99 62.20 70.00 -
Ours 68.24 77.90 67.45 77.19 83.35 48.84 37.20 87.83 70.04 87.12 62.29 69.77 -0.33%
Phi-3-mini-4k-instruct FP16 68.20 68.14 60.64 73.88 80.14 39.53 38.40 86.24 78.34 83.42 55.63 66.93
Ours 67.28 66.78 59.69 73.17 79.43 37.58 39.00 85.81 77.98 83.33 56.48 66.33 -0.90%
5 Conclusions and Limitations.
In this paper, we introduce SignRound, an efficient and succinct approach to optimizing weight rounding in quantization tasks. SignRound leverages signed gradient descent for lightweight block-wise tuning, resulting in significant quantization enhancements with minimal iteration steps. Our extensive experiments demonstrate that SignRound outperforms other quantization methods across various models and evaluation metrics while maintaining fixed hyperparameter configurations and similar fine-tuning time costs. Additionally, SignRound offers increased flexibility in quantizing specific large language models, achieving enhanced quantization performance with appropriately tuned hyperparameters.
Despite these significant advantages, we still observed a noticeable gap in the accuracy performance of ultra-low bit quantization compared to the original model, such as 2-bit quantization. This challenge could potentially be addressed by exploring non-uniform quantization and mixed-precision quantization, which we leave it as the future work.
Appendix A Quantiztion Cost
Table 8 compares the quantization costs of different methods, with all measurements conducted on a single NVIDIA A100 GPU with 80GB of memory. We ensure each evaluation process exclusively occupies one GPU, but CPU and other resources may be shared among different processes due to limited resource. For SignRound, we disabled low_gpu_mem_usage in our implementation to achieve faster tuning, albeit with higher memory usage. Nevertheless, LLaMAV2-70B was still able to run on an A100 GPU with 80GB of memory. Although HQQ is exceptionally fast, our methods outperform others in terms of speed. Table 9 also compares the costs between FlexRound, Adaptive Round, and our method.
Table 8: Quantization cost in seconds at W4G-1 for LLaMA2. Align with the accuracy experiments, Omniquant 70b is tested with 128 calibration samples, while all the others are tested with 512 samples.
Model GPTQ AWQ HQQ OmniQuant Ours
LLaMAV2-7B 1821 1328 19 10255 1041
LLaMAV2-13B 3266 2630 30 18186 1918
LLaMAV2-70B 18517 13586 119 35694 9116
Table 9: Quantization Time (seconds) of Rounding Methods at W4G-1 with 200 steps for LLaMA V2 Models and Mistral-7B.
Method FlexRound AdativeRound Ours
Mistral-7B-V0.1 9369 9332 1045
LLaMAV2-7B 9628 9701 1041
LLaMAV2-13B 17583 17865 1918
Appendix B View of distribution of tuned parameters
Figure 2 illustrates the distribution of the magnitudes of V in Eq.3 and $\alpha$,$\beta$ in Eq. 4 for Mistral-7B-v0.1 and LLaMA-2-7B at W4G-1. The results indicate that the distribution is flat for most layers, except for a few layers at the beginning and the end.
Mistral-7B, alpha values
Llama-2-7B, alpha values
Mistral-7B, beta values
Llama-2-7B, beta values
Mistral-7B, V values
Llama-2-7B, V values
Figure 2: The distribution of the magnitude of V in Eq. 3 and $\alpha$, $\beta$ in Eq. 4 for Mistral-7B-v0.1 and LLaMA-2-7B at W4G-1, each color in the distribution represents a specific layer index in the models, with blue indicating shallow layers closer to the data layer, and red representing deeper layers.
Table 10: Accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G-1. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues.
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
RTN 31.34 70.02 55.35 69.77 77.69 20.32 32.60 73.43 59.57 74.45 41.30 55.08
GPTQ 29.06 71.08 55.11 70.01 77.37 20.93 32.20 72.69 63.90 74.66 41.64 55.33
AWQ 33.33 70.81 55.98 68.27 78.07 21.18 31.40 74.37 64.62 74.03 41.21 55.75
Omni 32.52 72.13 55.87 70.17 78.35 22.77 32.80 75.05 66.07 75.13 40.19 56.46
Ours 31.80 71.96 56.57 69.53 79.00 21.91 33.20 75.72 66.79 74.83 43.09 56.76
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
RTN 39.57 70.93 58.82 71.98 78.02 24.85 32.00 78.20 66.43 75.67 44.62 58.28
GPTQ+ 40.01 74.67 58.92 71.03 78.45 26.44 33.60 77.09 68.23 76.85 44.97 59.12
AWQ 44.56 74.13 59.13 71.27 78.94 25.83 33.20 76.42 66.06 76.89 46.67 59.37
Omni 43.66 75.59 59.36 72.38 78.89 25.34 32.20 75.99 69.68 77.10 45.65 59.62
Ours 43.94 75.82 59.51 72.22 78.78 25.70 32.80 77.34 67.51 76.47 46.67 59.71
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
RTN 53.05 75.65 62.08 74.82 80.09 25.95 35.80 81.87 63.54 79.76 50.26 62.08
GPTQ 53.04 77.22 61.95 73.80 80.69 27.29 34.60 81.07 66.06 78.79 49.15 62.15
AWQ 54.13 76.77 62.78 74.11 81.07 27.78 35.00 82.66 67.15 79.97 51.71 63.01
Omni 53.43 77.64 62.73 75.30 80.58 26.56 35.40 82.51 67.87 79.76 50.51 62.93
Ours 54.72 77.84 62.91 75.06 80.69 26.68 36.40 82.60 66.79 80.13 52.13 63.27
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
RTN 58.74 76.42 64.12 76.72 81.01 29.25 38.60 84.13 70.40 80.72 51.88 64.73
GPTQ+ 59.10 78.17 63.78 75.69 81.34 28.27 38.40 83.76 68.59 80.98 51.62 64.52
AWQ 58.86 77.37 63.86 76.56 80.85 28.27 35.20 83.94 71.48 78.75 50.94 64.19
Omni 59.59 79.16 64.03 75.93 81.99 27.05 36.80 84.65 71.48 80.98 51.79 64.86
Ours 59.21 79.16 64.37 76.64 81.34 26.81 37.80 84.40 69.68 80.98 51.79 64.74
Table 11: Accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W4G128. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues.
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
RTN 59.72 74.44 61.06 73.40 80.36 27.17 32.60 83.67 64.62 79.63 49.32 62.36
GPTQ 59.17 74.52 60.37 74.90 80.58 26.68 31.00 83.33 67.15 79.67 48.12 62.32
Mistral-7B AWQ 60.20 75.14 60.43 73.80 80.03 27.05 30.40 84.01 62.09 80.39 50.26 62.16
HQQ 60.02 75.41 60.79 74.11 81.01 27.29 32.60 82.97 66.79 79.92 49.32 62.75
Omni 59.71 73.94 60.62 73.56 80.36 26.68 30.80 83.58 65.70 80.01 49.06 62.18
Ours 60.47 75.59 61.03 73.88 80.09 27.54 31.60 83.09 66.07 79.97 49.49 62.62
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
RTN 40.91 72.44 56.91 68.35 77.58 24.97 31.20 77.61 56.32 76.26 43.52 56.92
GPTQ 42.57 73.28 56.36 69.06 78.02 25.34 30.20 75.72 57.04 75.63 42.15 56.85
V2-7B AWQ 41.00 72.60 56.40 68.98 77.31 25.70 31.60 78.75 58.48 76.14 43.86 57.35
HQQ 41.79 73.20 56.21 68.43 77.58 25.83 31.60 76.09 62.82 75.84 42.15 57.41
Omni 41.72 73.04 56.59 68.98 77.91 24.97 30.80 75.81 61.37 75.76 43.34 57.30
Ours 41.82 72.75 56.79 68.67 78.13 25.58 30.20 77.49 63.54 75.76 42.58 57.57
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
RTN 52.10 76.27 59.77 72.14 78.62 24.72 34.20 80.24 62.09 79.00 47.95 60.65
GPTQ 52.66 76.54 59.76 72.14 78.35 25.70 34.00 79.33 66.43 78.58 47.53 61.00
V2-13B AWQ 52.39 76.89 59.97 73.24 79.00 25.21 32.60 80.40 63.54 79.04 47.70 60.91
HQQ 52.09 75.74 59.46 72.14 78.45 24.36 33.60 79.17 66.06 79.00 47.01 60.65
Omni 52.01 76.17 59.53 72.06 78.35 23.87 33.40 80.80 66.07 78.37 47.18 60.51
Ours 51.92 76.46 59.87 71.67 79.00 25.83 35.20 79.60 63.54 79.25 47.01 60.85
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
RTN 64.91 79.06 63.93 78.14 81.66 30.11 37.00 83.61 68.59 82.79 54.78 65.87
GPTQ 65.63 79.22 64.45 78.22 81.88 31.09 37.00 84.19 69.31 82.79 54.61 66.22
V2-70B AWQ 65.79 79.76 64.48 77.58 82.32 30.72 38.00 83.06 68.95 82.70 55.12 66.23
HQQ 65.34 79.14 64.56 77.35 81.56 30.48 37.20 83.67 69.31 82.83 55.20 66.06
Omni 65.30 79.39 64.52 77.51 81.88 30.60 37.40 83.39 68.23 82.91 55.12 66.02
Ours 65.65 79.49 64.60 78.30 82.05 31.58 37.40 84.83 68.95 82.87 54.52 66.39
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
RTN 32.63 72.31 56.26 70.01 78.45 20.93 33.60 74.74 64.26 74.71 42.75 56.42
GPTQ 31.16 72.40 55.85 70.09 78.13 22.28 30.40 74.65 64.26 74.20 40.19 55.78
AWQ 33.42 72.95 56.30 68.75 77.97 21.42 32.80 74.89 62.09 75.00 41.21 56.07
Omni 31.15 72.35 56.25 69.22 78.35 21.42 33.80 74.74 65.70 74.87 42.06 56.36
Ours 32.15 72.85 56.45 70.17 78.51 22.28 32.80 75.14 67.87 75.13 41.89 56.84
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
RTN 42.71 75.26 59.30 72.53 79.54 25.95 32.60 76.76 65.34 76.98 45.82 59.34
GPTQ+ 42.65 75.41 59.51 72.93 79.33 24.97 32.40 77.49 68.23 76.89 45.56 59.58
AWQ 42.66 75.76 59.50 72.77 78.89 26.56 33.60 77.46 68.59 76.94 45.48 59.84
Omni 43.99 76.29 59.53 73.56 79.43 25.83 33.20 77.58 67.15 76.64 45.48 59.88
Ours 42.27 76.17 59.53 73.56 79.33 25.70 32.80 78.20 70.04 76.94 46.25 60.07
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
RTN 54.24 77.02 62.90 74.35 80.52 27.29 34.20 81.96 67.15 80.89 52.05 62.96
GPTQ 54.20 77.41 62.79 75.14 80.41 27.54 34.60 81.93 67.51 80.05 50.51 62.92
AWQ 55.14 77.49 63.08 75.77 80.52 27.29 34.20 82.87 67.15 80.43 52.90 63.35
Omni 55.22 77.80 63.09 75.14 80.30 28.52 36.00 82.20 69.31 80.81 52.82 63.75
Ours 54.68 77.90 62.93 74.82 80.47 28.15 35.80 82.39 66.79 80.13 51.11 63.20
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
RTN 59.53 79.51 64.63 77.35 80.96 27.91 38.40 84.43 71.48 81.48 52.22 65.26
GPTQ+ 60.47 78.79 64.45 76.24 81.18 28.03 37.40 83.85 68.95 81.57 53.07 64.91
AWQ 59.45 79.31 64.67 76.72 81.56 28.15 38.00 84.43 71.12 81.10 52.13 65.15
Omni 59.27 78.65 64.48 76.87 81.23 27.78 39.00 84.13 70.76 81.57 53.07 65.17
Ours 58.93 79.22 64.48 77.03 81.28 27.91 38.60 84.31 70.76 81.19 52.22 65.08
Table 12: Accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W3G128. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues.
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
RTN 53.49 68.74 58.12 68.27 79.33 24.60 29.60 79.97 57.40 76.89 43.77 58.20
GPTQ 55.84 73.04 57.61 70.24 78.67 24.85 30.80 81.44 63.54 77.27 45.65 59.91
Mistral-7B AWQ 55.61 73.69 57.86 71.27 79.82 26.07 29.00 81.10 59.21 79.00 46.93 59.96
HQQ 53.97 68.66 58.59 72.22 78.73 25.70 30.00 80.24 63.90 76.81 43.86 59.33
Omni 54.79 69.34 58.42 68.51 79.38 24.85 28.80 80.15 56.68 77.74 45.14 58.53
Ours 57.54 73.01 59.60 72.85 79.54 25.70 31.60 81.74 58.12 78.70 46.33 60.43
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
RTN 34.22 65.96 54.90 67.56 76.28 24.48 30.80 71.68 54.51 72.98 38.57 53.81
GPTQ 36.11 69.61 53.66 68.59 76.01 21.91 27.80 73.43 54.51 73.74 40.19 54.14
V2-7B AWQ 35.82 69.90 54.98 67.40 76.01 25.21 29.80 74.68 57.76 74.07 41.64 55.21
HQQ 34.40 66.64 53.27 67.01 75.46 25.46 28.80 73.58 61.37 72.94 38.48 54.31
Omni 34.51 69.75 54.42 66.69 76.77 24.24 31.40 73.21 56.68 74.37 39.85 54.72
Ours 40.13 71.01 55.33 68.27 76.82 25.34 32.80 75.32 60.29 75.25 42.92 56.68
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
RTN 48.01 72.33 57.74 70.72 78.07 25.21 32.00 77.28 60.65 77.69 44.62 58.57
GPTQ 49.56 75.24 57.83 70.88 78.56 24.97 33.40 78.44 62.82 77.99 45.65 59.58
V2-13B AWQ 49.77 75.22 58.58 71.82 77.75 24.11 34.20 79.97 53.43 77.95 44.62 58.86
HQQ 48.40 73.22 57.66 69.77 77.31 24.11 30.60 76.97 60.29 77.15 43.60 58.10
Omni 47.25 73.67 58.46 70.01 78.40 24.36 33.60 79.79 64.62 77.86 46.16 59.18
Ours 49.64 75.20 59.11 71.59 78.29 24.85 34.20 78.47 58.12 78.58 45.82 59.44
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
RTN 61.15 77.95 61.98 77.90 80.79 29.74 36.00 81.28 64.62 81.10 52.39 64.08
GPTQ 63.15 79.06 62.94 77.66 81.45 30.72 36.20 81.53 67.87 81.65 53.67 65.08
V2-70B AWQ 64.09 79.47 63.75 76.48 81.77 29.74 37.20 82.69 66.06 81.40 53.67 65.12
HQQ 63.45 78.05 63.12 77.03 81.01 29.38 36.60 82.23 66.43 81.78 53.67 64.80
Omni 63.18 78.63 63.54 76.48 81.50 30.35 35.80 82.57 70.40 81.02 52.82 65.12
Ours 64.94 78.89 63.83 76.56 81.50 31.21 37.20 81.41 68.59 81.73 52.56 65.31
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
RTN 28.00 67.67 53.43 66.38 76.50 21.42 31.20 72.72 59.21 70.92 38.31 53.25
GPTQ 30.16 66.31 53.92 67.48 76.82 21.42 29.60 71.31 59.21 72.22 38.74 53.38
AWQ 30.33 70.19 54.53 68.98 76.71 20.81 31.60 74.68 64.62 73.23 38.91 54.96
Omni 28.35 70.54 54.48 68.27 77.48 21.05 29.40 72.29 66.07 72.73 37.12 54.34
Ours 25.85 70.95 55.45 69.69 77.37 21.66 32.00 73.88 60.29 73.48 39.33 54.54
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
RTN 34.87 69.65 57.25 70.48 77.31 26.93 32.00 71.44 62.82 75.63 43.94 56.57
GPTQ 35.51 73.08 57.89 70.80 77.37 24.48 31.40 77.52 62.82 74.41 43.26 57.14
AWQ 40.53 73.94 57.89 69.53 78.94 26.68 33.40 74.83 65.34 75.93 45.05 58.37
Omni 38.35 74.42 57.79 70.80 78.07 26.68 33.20 75.81 65.34 75.88 43.69 58.18
Ours 39.16 75.22 58.64 71.59 78.94 25.95 35.20 76.30 65.34 76.52 45.39 58.93
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
RTN 52.41 75.08 61.45 74.27 79.87 25.95 33.00 81.38 65.34 79.12 48.89 61.52
GPTQ 51.39 74.97 60.35 75.30 79.60 26.93 34.80 82.75 64.62 78.11 48.46 61.57
AWQ 53.84 76.71 61.94 75.14 80.03 25.34 34.40 81.90 67.15 79.59 50.77 62.44
Omni 53.67 76.95 61.82 74.51 80.14 25.95 34.40 81.10 66.07 79.76 48.21 62.05
Ours 54.39 77.49 62.13 74.03 80.47 27.30 35.00 79.76 68.59 79.46 48.98 62.51
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
RTN 57.47 77.43 63.23 75.93 80.41 28.64 38.40 82.69 66.43 80.22 51.19 63.82
GPTQ+ 57.92 78.69 62.98 76.87 80.63 27.66 37.60 84.16 68.95 80.89 51.19 64.32
AWQ 58.87 77.94 63.77 75.37 80.96 27.66 36.80 85.02 71.12 81.10 50.34 64.45
Omni 57.19 77.00 63.15 75.53 80.90 28.15 37.60 83.18 69.68 80.18 50.51 63.92
Ours 58.30 78.11 63.60 76.56 80.85 29.50 37.80 84.80 70.04 80.22 50.68 64.59
Table 13: Accuracies($\uparrow$) across 11 tasks(0-shot) of LLaMA and Mistral models at W2G128. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues.
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg.
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
RTN 23.45 0.14 27.43 49.64 54.30 24.24 15.20 38.69 51.99 29.08 21.59 30.52
GPTQ 25.23 30.47 38.28 53.83 64.91 24.11 17.40 58.29 50.90 47.77 24.57 39.61
Mistral-7B AWQ 25.38 0.00 25.71 52.01 51.58 23.99 17.60 37.83 47.29 26.98 22.27 30.06
HQQ 23.35 0.85 27.77 51.62 56.69 26.68 15.80 40.55 53.43 28.62 20.14 31.41
Omni 23.24 5.38 29.38 49.72 56.09 26.32 16.60 41.99 52.71 32.11 20.39 32.17
Ours 40.46 58.61 50.87 62.90 75.84 24.85 22.80 78.56 57.04 70.88 37.03 52.71
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
RTN 23.98 0.02 26.04 49.49 52.50 24.85 15.20 41.01 49.10 27.48 19.71 29.94
GPTQ 23.65 11.72 32.59 55.17 58.32 25.95 15.80 52.14 51.99 40.45 21.25 35.37
V2-7B AWQ 25.38 0.00 25.69 49.96 52.34 23.75 17.80 37.83 52.71 24.62 21.08 30.10
HQQ 24.51 0.02 26.06 49.49 53.26 24.72 13.80 37.92 50.90 26.52 21.33 29.87
Omni 22.97 35.53 40.28 55.88 65.13 22.89 15.60 63.24 53.07 50.13 23.46 40.74
Ours 27.20 55.25 47.35 61.01 72.96 24.85 25.60 68.07 54.51 65.99 32.25 48.64
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
RTN 23.77 7.47 33.08 49.01 57.94 26.19 16.00 47.74 53.43 32.03 21.93 33.51
GPTQ 24.69 45.20 41.06 55.80 67.08 23.26 19.80 54.40 52.35 55.60 27.82 42.46
V2-13B AWQ 27.04 0.00 25.80 51.85 52.99 23.62 13.60 62.17 47.29 26.22 23.12 32.16
HQQ 23.48 8.17 31.27 52.17 61.86 24.85 17.20 50.46 54.51 42.85 21.25 35.28
Omni 25.53 49.84 46.23 57.93 70.13 24.60 21.80 66.85 55.60 63.22 30.29 46.55
Ours 34.33 63.92 53.35 64.33 76.17 25.70 26.00 72.75 61.73 71.17 38.57 53.46
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
RTN 24.20 20.18 40.88 54.85 63.87 24.11 17.60 43.06 53.07 50.51 27.22 38.14
GPTQ 23.12 0.00 25.04 49.57 49.51 0.00 27.60 37.83 52.71 25.08 22.70 28.47
V2-70B AWQ 24.46 0.00 25.46 51.38 52.50 23.50 14.20 62.17 52.71 25.76 22.35 32.23
HQQ 23.16 19.46 35.45 56.67 66.00 22.52 20.00 40.46 52.71 52.06 23.12 37.42
Omni 33.84 61.83 52.44 64.33 74.10 24.48 28.20 71.68 53.07 67.21 33.28 51.31
Ours 54.04 72.97 59.65 74.90 79.00 29.01 34.80 79.63 69.68 78.37 46.59 61.69
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
RTN 24.36 0.52 27.24 49.25 54.24 24.24 15.20 39.63 57.40 27.86 21.84 31.07
GPTQ 22.95 12.75 33.36 51.70 60.07 23.99 13.40 48.62 53.07 40.82 21.50 34.75
AWQ 23.12 0.00 25.37 53.28 52.56 25.21 13.80 37.83 52.71 25.63 22.53 30.18
Omni 23.58 44.23 42.39 58.48 68.82 21.54 20.40 60.80 53.07 59.55 27.56 43.68
Ours 24.46 13.53 42.16 56.99 70.02 24.60 25.20 62.91 47.29 60.90 31.74 41.80
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
RTN 24.66 4.97 29.67 49.33 57.24 25.58 12.40 44.10 53.79 32.07 22.01 32.35
GPTQ+ 26.43 40.48 39.47 58.25 66.97 23.50 18.60 52.78 50.54 51.52 25.00 41.23
AWQ 27.04 0.00 25.59 50.36 53.05 24.11 15.60 62.17 47.29 25.97 23.21 32.22
Omni 26.93 56.41 47.67 61.17 73.23 23.38 24.60 68.75 53.07 67.00 33.79 48.73
Ours 31.87 59.65 51.25 67.64 76.28 25.58 27.80 69.11 58.48 70.71 37.12 52.32
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
RTN 23.24 5.55 27.22 53.99 56.80 21.79 18.20 51.65 53.07 36.74 21.33 33.60
GPTQ 30.47 49.93 45.05 61.88 68.88 23.26 22.60 68.29 51.99 60.69 30.72 46.70
AWQ 27.04 0.00 25.41 50.20 52.94 24.48 16.60 62.17 47.29 24.71 23.38 32.20
Omni 26.89 63.03 52.23 64.64 74.27 23.87 29.20 70.86 54.51 70.45 36.18 51.47
Ours 40.83 67.92 56.73 68.90 76.17 24.36 31.60 75.54 62.45 74.92 42.41 56.53
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
RTN 24.48 32.78 43.59 57.85 67.52 22.89 22.80 61.53 50.54 52.10 28.24 42.21
GPTQ+ 37.06 67.44 53.97 69.46 76.44 24.36 28.00 73.64 60.29 71.34 38.57 54.60
AWQ 25.38 0.00 25.58 49.96 53.10 24.24 11.00 37.83 52.71 24.96 22.44 29.75
Omni 27.36 65.94 55.53 68.11 76.99 25.21 29.60 75.69 59.21 69.82 35.07 53.50
Ours 47.21 72.07 60.06 73.24 78.62 25.46 34.20 80.64 62.82 77.48 46.76 59.87
Table 14: Perplexity(PPL) ($\downarrow$) of Wikitext2, PTB , C4 and Wikitext tasks for LLaMA and Mistral models. we follow the source code of GPTQ for wikitext2, PTB and C4 PPL evaluation, while for wikitext, we adopt lm-eval-harness [Gao et al., 2023]. NAN indicates no data acquisition, while OOM denotes out of memory.
LLaMA-V2 Wiki2. Ptb C4 Wiki.
7B FP16 5.47 37.92 7.26 8.79
W4G-1 RTN 6.12 82.85 8.16 10.06
GPTQ 5.84 1246 7.82 9.59
AWQ 5.81 57.09 7.70 9.42
Ours 7.85 3005.52 7.71 10.34
W4G128 RTN 5.72 65.35 7.58 9.22
GPTQ 5.60 246.28 7.48 9.05
AWQ 5.61 42.67 7.44 9.03
Ours 8.96 473.78 7.50 9.01
W3G128 RTN 6.66 55.10 8.98 11.21
GPTQ 6.32 2245 8.55 10.37
AWQ 6.24 66.57 8.27 10.18
Ours 8.09 164.90 8.12 9.76
W2G128 RTN 4270 9646 4807 1.8e5
GPTQ 25.56 9429 34.87 79.65
AWQ 2.3e5 2.1e5 1.7e5 1.1e7
Ours NAN NAN NAN NAN
13B FP16 4.88 50.93 6.73 7.90
W4G-1 RTN 5.20 60.69 7.14 8.65
GPTQ 5.12 55.99 7.04 942.3
AWQ 5.07 55.39 6.96 8.39
Ours 5.00 51.71 6.89 8.33
W4G128 RTN 4.98 53.69 6.87 8.12
GPTQ 4.98 52.43 6.85 10.86
AWQ 4.97 54.18 6.84 8.08
Ours 4.96 51.62 6.83 8.14
W3G128 RTN 5.52 64.85 7.58 9.27
GPTQ 5.39 72.96 7.47 334.2
AWQ 5.30 57.66 7.30 8.81
Ours 5.23 53.82 7.18 8.68
W2G128 RTN 122.5 1212 131.8 1054
GPTQ 11.30 410.9 15.11 270.6
AWQ 1.2e5 1.1e5 9.7e4 5.5e6
Ours 7.64 4250 11.73 57.52
70B FP16 3.32 24.25 5.71 4.54
W4G-1 RTN 3.67 23.56 6.01 5.18
GPTQ 3.57 23.76 5.89 5.00
AWQ 3.48 24.93 5.85 4.81
Ours 3.44 24.33 5.81 4.78
W4G128 RTN 3.46 24.20 5.83 4.78
GPTQ 3.42 24.01 5.78 4.71
AWQ 3.41 24.36 5.77 4.70
Ours 3.40 23.69 5.77 4.68
W3G128 RTN 3.98 23.59 6.27 5.77
GPTQ 3.83 24.78 6.09 5.50
AWQ 3.73 25.68 6.03 5.31
Ours 3.68 24.26 5.99 5.23
W2G128 RTN 27.01 758.9 47.57 298.3
GPTQ NAN NAN NAN NAN
AWQ 7.2e4 8.1e4 NAN 2.5e6
Ours NAN NAN NAN NAN
Mistral Wiki2. Ptb C4 Wiki.
7B FP16 5.25 35.00 8.38 OOM
W4G-1 RTN 5.99 44.88 9.47 OOM
GPTQ 5.57 54.45 8.86 OOM
AWQ 5.75 42.21 9.14 OOM
Ours 5.43 81.67 8.66 OOM
W4G128 RTN 5.42 34.08 8.62 OOM
GPTQ 5.37 37.53 8.56 OOM
AWQ 5.37 37.12 8.55 OOM
Ours 5.34 36.36 8.51 OOM
W3G128 RTN 6.16 49.97 9.68 OOM
GPTQ 5.90 49.50 9.30 OOM
AWQ 5.90 51.01 9.27 OOM
Ours 5.66 44.50 8.96 OOM
W2G128 RTN 1375 2351 1015 OOM
GPTQ 16.59 269.2 22.38 OOM
AWQ 3.7e4 3.4e4 3.7e4 OOM
Ours 8.70 86.08 12.54 OOM
LLaMA-V1 Wiki2. Ptb C4 Wiki.
7B FP16 5.68 41.15 7.34 9.49
W4G-1 RTN 6.29 48.65 8.12 10.62
GPTQ 6.13 47.18 7.93 10.32
AWQ 5.97 48.25 7.73 10.11
Ours 5.93 54.84 7.62 9.91
W4G128 RTN 5.96 42.33 7.70 10.00
GPTQ 5.90 42.36 7.66 9.91
AWQ 5.80 44.00 7.50 9.75
Ours 5.79 56.45 7.49 9.74
W3G128 RTN 7.01 56.28 9.18 12.11
GPTQ 6.60 53.75 8.72 11.46
AWQ 6.32 49.27 8.21 10.81
Ours 6.28 47.57 8.09 10.55
W2G128 RTN 1847 6574 936.2 1.3e4
GPTQ 28.52 638.3 37.85 128.0
AWQ 2.6e5 2.8e5 2.9e5 2.1e7
Ours 641.8 824.9 2533 1876
13B FP16 5.09 28.10 6.80 14.06
W4G-1 RTN 5.53 29.45 7.23 37.17
GPTQ 5.34 30.23 7.09 13.09
AWQ 5.25 30.34 7.01 12.36
Ours 5.21 27.81 6.93 113.24
W4G128 RTN 5.26 28.36 6.94 25.34
GPTQ 5.19 29.36 6.91 13.33
AWQ 5.19 28.34 6.90 15.25
Ours 5.18 27.80 6.88 59.09
W3G128 RTN 5.88 33.10 7.86 44.06
GPTQ 5.56 32.52 7.48 95.24
AWQ 5.53 29.63 7.34 22.26
Ours 5.45 28.13 7.21 15.44
W2G128 RTN 797.7 1695 449.1 1.5e4
GPTQ 12.13 185.8 NAN 546.1
AWQ 2.8e5 2.6e5 2.4e5 1.6e7
Ours 8.36 48.93 10.64 1773
30B FP16 4.10 23.51 6.13 6.89
W4G-1 RTN 4.54 25.49 6.54 8.03
GPTQ 4.41 24.22 6.40 8.50
AWQ 4.30 24.20 6.30 6.88
Ours 4.23 27.97 6.24 6.90
W4G128 RTN 4.23 23.90 6.26 7.05
GPTQ 4.24 23.92 6.23 7.73
AWQ 4.22 23.98 6.21 7.29
Ours 4.18 31.38 6.20 7.39
W3G128 RTN 4.87 26.99 6.85 NAN
GPTQ 4.72 25.14 6.73 8.44
AWQ 4.61 25.05 6.56 7.84
Ours 4.50 67.01 6.47 7.90
W2G128 RTN 68.40 566.8 114.2 1192
GPTQ 9.21 59.75 12.50 21.21
AWQ 2.3e5 2.2e5 2.4e5 1.5e7
Ours 7.13 55.40 12.02 118.7
LLaMA-V1 Wiki2. Ptb C4 Wiki.
65B FP16 3.53 25.07 5.81 4.96
W4G-1 RTN 3.92 28.07 6.07 5.60
GPTQ 3.79 34.82 6.00 5.46
AWQ 3.72 44.83 5.96 5.30
Ours 3.65 22.42 5.89 5.19
W4G128 RTN 3.67 25.61 5.90 5.21
GPTQ 3.64 33.81 5.88 5.17
AWQ 3.62 24.46 5.87 5.14
Ours 3.61 35.87 5.87 5.13
W3G128 RTN 4.25 50.00 6.33 6.25
GPTQ 4.05 32.64 6.21 6.03
AWQ 3.95 23.48 6.14 5.83
Ours 3.90 29.15 6.08 5.69
W2G128 RTN 15.21 276.7 20.03 29.39
GPTQ 6.85 37.79 NAN 12.25
AWQ 7.3e4 6.7e4 7.4e4 NAN
Ours 5.52 NAN NAN 9.25
Appendix C Other results
We present the detailed 11 tasks accuracy results for llama and mistral models with a size of 7-70B at W2-W4 in Table 10, 11, 12 and 13. the detailed PPL results in Table 14. Overall, our SignRound demonstrates a clear advantage in accuracy tasks, particularly in ultra-low bit quantization, achieving state-of-the-art performance compared to several popular weight quantization methods. In terms of perplexity (PPL), SignRound outperformed all other methods in 57 out of 112 scenarios, demonstrating its advantages. However, we observed that several quantization algorithms, including SignRound, exhibit sensitivity across different models and tasks. The reason for this sensitivity is detailed in Section 4.1.